Ask your own question, for FREE!
Mathematics 9 Online
OpenStudy (anonymous):

CAN SOMEONE PLEASE HELP ME!!!! 20. A drawer contains 4 red socks, 3 white socks, and 3 blue socks. Without looking, you select a sock at random, replace it, and select a second sock at random. What is the probability that the first sock is blue and the second sock is red? (4 points)

OpenStudy (anonymous):

PLEASE HELP REALLY NEED THIS

OpenStudy (cwrw238):

P(first blue) = 3/10 P(red second) = 4/10 as these actions are independent of each other the probabilities are multiplied so required probability is 3/10 * 4/10 = 12/100 = 3/25

OpenStudy (anonymous):

thanks are you sure this is correct it could make or break my grrade

OpenStudy (anonymous):

one question, this question, cannot make/break your grade... :)

OpenStudy (anonymous):

if its worth four points and its you last test for the school year

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
1 day ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
1 day ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!